OpenStudy (anonymous):

42w>2(w+7) im confused

OpenStudy (anonymous):

21w > w+7 20w > 7 w > 7/20

OpenStudy (ash2326):

@mimimarie416 do you get this?

OpenStudy (anonymous):

not really

OpenStudy (ash2326):

we have \[42w>2(w+7)\] Divide both sides by 2 \[\frac{42w}2>\frac{2(w+7)}{2}\] \[\frac{\cancel{42}21w}{\cancel 2}>\frac{\cancel2(w+7)}{\cancel 2}\] we get \[21w>w+7\] do you get this part?